OpenStudy (anonymous):

Find the equation of the line that contains P (-1,3) and is perpendicular to the line that contains A (-2,0) and B (4,3)

OpenStudy (wilsondanielle):

Start by using the points of the second line to find its slope. Use that slope and point P to find the intercept. This will yield the full equation of the line.

OpenStudy (wilsondanielle):

Slope = rise/run = \[\Delta y / \Delta x\] = 3/6 = 0.5 3 = 0.5 (-1) + intercept 3=-0.5 + intercept intercept = 3+0.5 = 3.5 therefore the equation would be y=0.5x+3.5